# An unidentified Fermi source emitting radio bursts in the Galactic bulge

### v1

@@ 3 Periodicity Search
Given the millisecond-durations of the bursts, the simplest explanation of this source is that it is a neutron-star related phenomenon. Thus, we aim to determine whether there is any periodicity in the pulse arrival times. We carried out several periodicity search techniques.
@@ 3.1 Fast-Fourier-transform (FFT) searches
We carried out FFT searches using the standard pulsar analysis package PRESTO 333https://github.com/scottransom/presto. As per standard PRESTO procedure, we mitigated RFI using the rfifind mask, and created dedispersed time series of the data using prepdata in the DM ranges 995-1025 pc cm. We ran FFT on the resulting time series to look for periodic signals using the code realfft, followed by accelsearch, which runs Fourier-domain acceleration searches and harmonic summing, with zmax=200. We also ran the search without harmonic summing (zmax=0). We then folded the time-series data at all candidate periods with a S/N using the function prepfold. There were multiple S/N candidates in all searches. Later, the frequency-time filterbank files at each epoch were dedispersed at their respective average single-pulse DM (: Epoch 1=1010 pc cm, Epoch 2L =1012 pc cmand Epoch 2C=1001 pc cm), and folded at all the periods reported by the accelsearch above S/N. However, when the plots were visually inspected, it was apparent that all of these significant candidates were due to isolated instances of narrow-band RFI. The search was done in the barycenter frame of reference.
@@ 3.2 Fast Folding Analysis
We also carried out a fast-folding analysis (“FFA”), which allows superior sensitivity to long-period (1 s) candidates than the FFT search. The minimum separation two bursts (L3 and L4), 7.09 minutes, sets the upper limit to the possible period. Therefore we use 8 min (480 s) as the maximum period we search for in our trials. We ran the FFA software riptide (Morello et al., 2020) on the time series data produced from presto for all three epochs, dedispersed at the average DM at the respective epoch. We ran three separate searches on the time ranges from 0.01 - 1 s, 1 - 10 s and 10 - 480 s, analyzed the periodogram and folded the time series at the highest S/N period detection. The folded sub-integrations plot had no significant detections in all three epochs.
@@ 4 Burst properties
where , , is the area, standard deviation, and mean of the Gaussian pulse, and is the scattering tail. Some of the bursts were fitted with multiple components, if adding components resulted in a reduced chi-squared value closer to 1. The MJDs reported are derived from the of the pulse and are referenced to infinite frequency and to the barycentre. We didn’t have a flux calibrator for epoch 1 L-band bursts (L1–L4). Therefore, for that epoch only, the flux density is derived from the radiometer equation given by:
@@
where and is the off-pulse standard deviation and mean. is the equivalent width of the pulse in bins. This flux density was multiplied by the width of the burst to get the fluence.
@@ 4.1 Polarimetry
Bursts in epoch 1 did not have a polarization calibrator available. For all epoch 2 bursts (L and C), we calibrated polarization using the standard package pac. We did 1-D RM Synthesis (Brentjens & de Bruyn, 2005; Heald, 2009) on the Stokes data to search an RM range between rad mand found an average RM of 816 rad m with a standard deviation of 9.82 rad m. We corrected the Faraday rotation by de-rotating the bursts at their respective RMs using the psrchive package pam. No RM was detected for burst C3. The bursts were averaged in frequency, and the unbiased linear polarization was calculated using the following (Everett & Weisberg, 2001):
@@ 4.2 Spectral index and Energetics
The detection of bursts at different frequencies ranging from 1.1–1.9 MHz and 4–8 GHz allows us to calculate the spectral index for the source. Due to the large bandwidth, we split the C-Band burst flux into three equal subbands to allow a better spectral quantification. All the ten bursts from Epoch 1 and Epoch 2L and 2C are plotted in Figure. 4. The flux of each burst was calculated by dividing the total fluence of each burst by the total width. The error on the flux is calculated by propagating the errors on fluence and width. Assuming a power law variation of flux with frequency, , we fitted a curve with two free parameters given by:
where is the y-axis intercept and is the power-law index, using scipy.optimize.curvefit. The error bars on the fluxes were accounted for while fitting the curve. Here, the slope of the curve in the log scale is equivalent to the spectral index of the source. The source has a steep spectral index of -2.70.03. We also calculated the in-band spectral index between 4 and 8 GHz using the brightest C-Band burst C1 and measured .
Using the DM-distance, we also calculated the specific luminosity of the source at 1.4 GHz assuming an opening angle of 6 which is given by:
The maximum specific luminosity of all the bursts in the full-pulse sample is () using d=10 kpc (5 kpc), from burst L4. We also calculated the pseudo luminosity given by for the full pulse and sub-pulse sample is plotted in Fig. 5 with other fast radio transient sources. We can calculate the brightness temperature of the source using
For a peak flux of 0.32 Jy and a width of 14.6 ms at 1.4 GHz, we get the brightness temperature to be about and this implies a coherent emission mechanism.
@@ 4.3 Scattering
The scattering timescale was measured by assuming that the burst profile was a Gaussian convolved with a one-sided exponential tail with a delay given by . The bursts in our sample show scattering at both L and C bands, as reported in Table 2. There is a noticeable difference in the scattering timescale between the two epochs of 1.4 GHz observations with a mean value of 26.561.40 ms in epoch 1 and 17.630.69 ms in epoch 2. This implies a variable scattering screen between the source and the observer. The scattering timescale reported for C-Band bursts could be mischaracterization due to complex burst morphology.
@@ 5 Discussion
First, we present a summary of basic observations. The location of the source towards the Galactic plane explains the high DM, and the source is decidedly of Galactic origin. The smallest burst duration of 1.6 ms suggests that the object may have a radius of a few hundred kilometers. The high brightness temperature of K suggests a coherent emission mechanism. The high fraction of linear polarization implies ordered magnetic fields. The bursts show stochastic variations in DM around a mean of 1009 pc cmand a standard deviation of 6 pc cm. The scattering timescale decreased by about 40% between the first and second epoch L-Band observation, suggesting a dynamic scattering screen. This variation appears to be uncorrelated with DM variations. The relatively low luminosity and, accordingly, low brightness temperatures argue away from a Galactic-FRB scenario for the pulses.
@@
We examined other observations of the field to determine whether any objects might be colocated or associated with the bursts at the realfast position. Figure 6 summarizes the multi-frequency emission within approximately 0.25 of the bursts. Ultimately, we come to the conclusion that while there is a star within the positional localization error region, there are no confidently associated stars, pulsars, or supernova remnants with the burst source. The Fermi source, however, does appear to be confidently associated with the bursts.
@@ 5.1.1 Co-located stars and pulsars
There are four known pulsars within 0.25 degrees, as listed in the ATNF Pulsar Database Manchester et al. (2005), however these appear to be unrelated given their much different positions and DMs ( pc cm). The nearest of these is still approximately 10 arcminutes away from the realfast position, thus, none are spatially coincident.
A search of SIMBAD (Wenger et al., 2000) and the NASA Extragalactic Database returned no objects within three times the RMS realfast position error (RA and Dec of 0.22, ); the nearest catalogued object was what appears to be an unrelated 2MASS source 5 to the south-east. To check for any faint uncataloged stars colocated with the bursts, we searched the PAN-STARRS DR2 catalog within 2 radius. This returned one object ID within this region. The object, PSO J274.6436-15.5258 (object ID 89362746437589401) has a reported DR2 stacked-image position (J2000 , ), and is the bright compact source 1.58 north-west from the realfast position shown on the inset of Fig. 6. This source is consistent with the realfast position within the error region quoted on the Pan-STARRS source.
Using the classification scheme set by (Tachibana & Miller, 2018), we obtain a crude object type identification (star vs. galaxy) by comparing the difference between the reported i-band PSFMag and KronMag values (computed as iMeanPSFMag–iMeanKronMag) in the PAN-STARRS catalog. The iMeanPSFMag value of 20.283 and difference value of –0.07 imply that this object is more likely to be a star than a galaxy. Using the iKronMag, zKronMag and the yKronMag values, we computed the i–z difference and the z–y difference. The differences were 0.91 and 0.47, respectively, which makes it consistent with the star having the spectral classification M5V and mass 0.16 (Pecaut & Mamajek, 2013). Given the proximity of this field to the Galactic center, it is expected that some spatially coincident emissions may arise by chance. We determined an empirical estimate for chance coincidence of sky position overlap by noting that there are 77 sources in the PAN-STARRS catalog within a 20 radius of the realfast position; thus the source density in this image is approximately 794119 deg. The chance coincidence of our search finding a source in a radius of 1.58 is ; this neither rules out nor supports an association between the bursts and the candidate star. That is, the density of stars in the field is too high to associate the objects confidently by spatial coincidence.
We also queried NASA’s The High Energy Astrophysics Science Archive Research Center (HEASARC) within a radius of 2 to search for any high energy counterpart. The query returned an XMM Newton Source at J2000 RA and Dec 18:18:34.639, –15:31:35.1 with an error radius of 1.316. It has a combined EPIC band 8 flux of . For a distance of 10.9 kpc, this would become an X-ray luminosity of . The position of this X-ray source is not consistent with the realfast position within the formal error, but is separated at a distance within 2 of the positional error.
@@ 5.1.2 Supernova Remnant and Pulsar Wind Nebula
It does not appear that our burst source is a pulsar associated with this PWN or supernova event, based simply on its large offset; if it is associated with the PWN, the pulsar (or Fermi emission) should be directly coincident with that emission. Neither the Fermi or realfast error regions are coincident with the PWN emission. The realfast position sits at a distance of 6.1 from the central position of the PWN. At the distance of 9.3 kpc indicated by Su et al. (2017), and their modelled supernova age of 11 kyr, the pulsar would have had to travel at a velocity of 1400 km s to reach its current position (this minimum assumes the trajectory is perfectly orthogonal to our line-of-sight). This is much larger than typical pulsar birth velocities in the Milky Way (Lorimer et al., 1997). While there is some precedent for such a configuration (Schinzel et al., 2019), the lack of morphological evidence of such a high velocity (trail or bow shock) in the VLA and archival data, in addition to the central PWN, argue against that scenario.
@@ 5.1.3 Statistical association of the Fermi error regions
First, it is clear that the Fermi 3FGL position is slightly different from that of 4FGL. Both 95% Fermi error regions are consistent with the realfast bursts. The 4FGL position has a much smaller error region, and we characterize a conservative likelihood of chance coincidence for this occurring by answering the question, “given a random realfast position, what’s the chance it would have landed within a random 4FGL error region of this size?” We do this by estimating the density of Fermi sources (number of sources per square degree) near the Galactic center region, then multiplying this by the angular coverage of the 4FGL region in Figure 6. In the 4FGL catalog, we find 559 sources within the region covering in Galactic longitude and in Galactic latitude around the Galactic center. This tells us the source density per sky area is around 0.7 sources per square degree. Multiplying that by the area of the 4FGL ellipse (minor axis 144.36, major axis 194.76), we find a probability of 0.0048 that the realfast position would overlap the 4FGL error region by chance. This number would be less if we considered only unidentified Fermi sources, and took into account the additional confidence afforded by the 3FGL error region; thus, we consider the colocation, and likely association of the Fermi source with this burst emitter, to be relatively secure.
If the source were associated with a -ray pulsar then following equation 24 of the Smith et al. (2023) using the DM distance and a G100 = erg cm s, which yields a -ray luminosity of erg s with . If the nature of the detected pulses were from a pulsar-like object, this would put it into the category of young pulsars.
@@ 5.2.1 Magnetars
Magnetars are a class of rotating neutron stars with very high magnetic fields, and belong to the younger class of neutron stars. They can exhibit a rotation period of 1-10 s and very high spin down up to . They typically show time-variable flux densities, high linear polarization, profile changes, and flat spectral index (Levin et al., 2012). Some magnetars are known to show an inverted spectrum (Camilo et al., 2007a; Levin et al., 2012). Some magnetars also show circular polarization (Levin et al., 2012; Kramer et al., 2007). These are also known -ray and X-ray emitters in keV and MeV ranges. Their decaying magnetic field powers their emission, and they are capable of emitting radio bursts of energies similar to extragalactic FRBs, up to ergs(Bochenek et al., 2020; CHIME/FRB Collaboration et al., 2020).
The high polarization fraction, variable fluxes, and energies of J1818–1531 are consistent with a magnetar origin. A similar steep spectral index has also been observed in J1818–1607 (Lower et al., 2020). Magnetars are also known to have a large duty cycle up to 50% (Levin et al., 2012), which can lead to a non-detection of a period. The gradual change or flatness of PA is also commonly seen in magnetars (Camilo et al., 2007b).
Magnetars have not been detected in GeV energies, except for one extragalactic magnetar giant flare detected by Fermi-LAT (Fermi-LAT Collaboration et al., 2021). However, that flare was transient (fading on a timescale of less than 300 seconds). The high energy magnetar thermal emission typically arises in the X-ray band and is modelled by a blackbody in the soft X-ray band and a power law component in the hard X-rays. Kong et. al predicts non-thermal -ray luminosity from magnetars of the order in the GeV energies and suggests that the GeV photons can only escape the magnetar surface if the corresponding X-ray luminosity is . If the XMM Newton source is associated with J1818–1531, then the X-ray and -ray energies are consistent with this model and this will be the first detection of GeV energies from a Galactic magnetar.

### v2

@@ 3 Periodicity Search
Given the millisecond-duration of the bursts, the simplest explanation of this source is that it is a neutron-star related phenomenon. Thus, we aim to determine whether there is any periodicity in the pulse arrival times. We carried out several periodicity search techniques.
@@ 3.1 Fast-Fourier-transform (FFT) searches
We carried out FFT searches using the standard pulsar analysis package PRESTO 333https://github.com/scottransom/presto. As per standard PRESTO procedure, we mitigated RFI using the rfifind mask, and created dedispersed time series of the data using prepdata in the DM ranges 995-1025 pc cm-3. We ran FFT on the resulting time series to look for periodic signals using the code realfft, followed by accelsearch, which runs Fourier-domain acceleration searches and harmonic summing, with zmax=200. We also ran the search without acceleration (zmax=0). We then folded the time-series data at all candidate periods with a S/N using the function prepfold. There were multiple S/N candidates in all searches. Later, the frequency-time Filterbank files at each epoch were dedispersed at their respective average single-pulse DM (: Epoch 1=1010 pc cm-3, Epoch 2L =1012 pc cm-3and Epoch 2C=1001 pc cm-3), and folded at all the periods reported by the accelsearch above S/N. However, when the plots were visually inspected, it was apparent that all of these significant candidates were due to isolated instances of narrow-band RFI. The search was done in the barycenter frame of reference.
@@ 3.2 Fast Folding Analysis
We also carried out a fast-folding analysis (“FFA”), which allows superior sensitivity to long-period (1 s) candidates than the FFT search. The minimum separation between two bursts (L3 and L4), 7.09 minutes, sets the upper limit to the possible period. Therefore we use 8 min (480 s) as the maximum period we search for in our trials. We ran the FFA software riptide (Morello et al., 2020) on the time series data produced from presto for all three epochs, dedispersed at the average DM at the respective epoch. We ran three separate searches on the time ranges from 0.01 - 1 s, 1 - 10 s and 10 - 480 s, analyzed the periodogram and folded the time series at the highest S/N period detection. The folded sub-integrations plot had no significant detections in all three epochs.
@@ 4 Burst properties
where , , is the area, standard deviation, and mean of the Gaussian pulse, and is the scattering tail. Some of the bursts were fitted with multiple components, if adding components resulted in a reduced chi-squared value closer to 1. The frequency channels that were zapped were removed before averaging. The MJDs reported are derived from the of the pulse and are referenced to infinite frequency and to the barycentre. We didn’t have a flux calibrator for epoch 1 L-band bursts (L1–L4). Therefore, for that epoch only, the flux density is derived from the radiometer equation given by:
@@
where and is the off-pulse standard deviation and mean. This flux density was multiplied by the width of the burst (total width in case of L2) obtained from profile fitting to get the fluence.
@@ 4.1 Polarimetry
Bursts in epoch 1 did not have a polarization calibrator available. For all epoch 2 bursts (L and C), we calibrated polarization using the standard package pac. We did 1-D RM Synthesis (Brentjens & de Bruyn, 2005; Heald, 2009) on the Stokes data to search an RM range between rad m-2and found an average RM of 816 rad m-2 with a standard deviation of 9.82 rad m-2. We corrected the Faraday rotation by de-rotating the bursts at their respective RMs using the psrchive package pam. There were no significant peaks in the Faraday Dispersion Function of C3. Therefore, we de-rotated C3 at the mean RM of C1 and C2 i.e 828 rad m-2. The bursts were averaged in frequency, and the unbiased linear polarization was calculated using the following (Everett & Weisberg, 2001):
@@ 4.2 Spectral index and Energetics
The detection of bursts at different frequencies ranging from 1.1–1.9 MHz and 4–8 GHz allows us to calculate the spectral index for the source. Since the burst detections are non contemporaneous and has a large gap between them, we have to make the assumption that the source is stable during this timescale and does not show any intrinsic flux variability at a given frequency. All the 15 bursts from the VLA/realfastand GBT Epoch 1 and Epoch 2L and 2C are plotted in Figure. 5. The 8 value of VLA S-band is also included for completeness. The flux of each burst in epoch 2L and epoch 2C, was calculated by dividing the total fluence of each burst by the total width. The error on the flux is calculated by propagating the errors on fluence and width. The flux of Epoch-1 L-band bursts are calculated as described in §4 and the error bar is calculated from the radiometer noise. Assuming a power law variation of flux with frequency, , we fitted a curve with two free parameters given by:
where is the y-axis intercept and is the power-law index, using scipy.optimize.curvefit. The error bars on the fluxes were accounted for while fitting the curve. Here, the slope of the curve in the log scale is equivalent to the spectral index of the source. The source has a spectral index of –1.50.036.
Using the DM-distance, we also calculated the isotropic luminosity of the source at 1.4 GHz given by . The maximum isotropic luminosity of all the bursts in the full-pulse sample is () using d=10.9 kpc (5.8 kpc), from burst L6. We also calculated the pseudo luminosity given by for the full pulse and sub-pulse sample is plotted in Fig. 6 with other fast radio transient sources. We can calculate the brightness temperature of the source using
The brightness temperature distribution for all the bursts components has a mean value of and a 1 value of for NE2001(YMW16) distances. Note that we removed burst L2 from this calculation to avoid any inconsistencies.
@@ 4.3 Scattering
The scattering timescale was measured by assuming that the burst profile was a Gaussian convolved with a one-sided exponential tail with a delay given by . The bursts in our sample show scattering at both L and C bands, as reported in Table 2. There is a noticeable difference in the scattering timescale between the two epochs of 1.4 GHz observations with a mean value of 26.561.40 ms in epoch 1 and 17.630.69 ms in epoch 2. This implies a variable scattering screen between the source and the observer. The scattering timescale reported for C-Band bursts could be mischaracterization due to complex burst morphology. To test the credibility of the scattering timescale variation between epochs, we plotted the scattering timescale vs the frequency for the bursts L4 from Epoch 1 and burst L6 from Epoch 2. The bursts were selected based on the criteria that they have enough S/N to be split into different bands and they have good model fit. The sub-bands with poor fit were removed. Three sub-bands of L4 and 6 sub-bands of L6 are used. The data was fit with . The best fit curve for L4 has a and L6 has a (Fig. 7). For Kolmogorov spectrum, . The index of scattering timescales in both epochs are consistent within error bars with that of a Kolmogorov medium and therefore suggests that the epoch-to-epoch variability seen in L-band bursts are indeed real.
@@ 5 Discussion
First, we present a summary of basic observations. The location of the source towards the Galactic plane explains the high DM, and the source is decidedly of Galactic origin. The smallest burst duration of 1.6 ms suggests that the object may have a radius of a few hundred kilometers. The high brightness temperature of K suggests a coherent emission mechanism. The high fraction of linear polarization implies ordered magnetic fields. The bursts show stochastic variations in DM around a mean of 1009 pc cm-3and a standard deviation of 6 pc cm-3. The scattering timescale decreased by about 40% between the first and second epoch L-Band observation, suggesting a dynamic scattering screen. This variation appears to be uncorrelated with DM variations. The relatively low luminosity and, accordingly, low brightness temperatures argue away from a Galactic-FRB scenario for the pulses for now, even though more energetic bursts similar to the ones seen from SGR 1935+2154 cannot be excluded to occur in the future.
@@
We examined other observations of the field to determine whether any objects might be co-located or associated with the bursts at the realfast position. Figure 8 summarizes the multi-frequency emission within approximately 0.25∘ of the bursts. Ultimately, we come to the conclusion that while there is a star near the positional localization error region, there are no confidently associated stars, pulsars, or supernova remnants with the burst source. The Fermi source, however, does appear to be confidently associated with the bursts.
@@ 5.1.1 Co-located stars and pulsars
There are four known pulsars within 0.25 degrees, as listed in the ATNF Pulsar Database (Manchester et al., 2005), however these appear to be unrelated given their much different positions and DMs ( pc cm-3). The nearest of these is still approximately 10 arcminutes away from the realfast position, thus, none are spatially coincident.
A search of SIMBAD (Wenger et al., 2000) and the NASA Extragalactic Database returned no objects within three times the RMS realfast position error (RA and Dec of 0.22′′, ); the nearest cataloged object was what appears to be an unrelated 2MASS source 5′′ to the south-east. To check for any faint uncataloged stars co-located with the bursts, we searched the PAN-STARRS DR2 catalog within a 2′′ radius. This returned one object ID within this region. The object, PSO J274.6436-15.5258 (object ID 89362746437589401) has a reported DR2 stacked-image position (J2000 , ), and is the bright compact source 1.58′′ north-west from the realfast position shown on the inset of Fig. 8. This source is consistent with the realfast position within the error region quoted on the Pan-STARRS source. The source is identified by the GAIA catalogue as being of a Star object type with a 99.95% probability. This is consistent with the PAN-STARRS measurements using the classification scheme set by (Tachibana & Miller, 2018), where we can obtain a crude object type identification (star vs. galaxy) by comparing the difference between the reported i-band PSFMag and KronMag values (computed as iMeanPSFMag–iMeanKronMag) in the PAN-STARRS catalog. The iMeanPSFMag value of 20.283 and difference value of –0.07 imply that this object is more likely to be a star than a galaxy. Using the iKronMag, zKronMag and the yKronMag values, we computed the i–z difference and the z–y difference. The differences were 0.91 and 0.47, respectively, which makes it consistent with the star having the spectral classification M5V and mass 0.16 (Pecaut & Mamajek, 2013). Given the proximity of this field to the Galactic center, it is expected that some spatially coincident emissions may arise by chance. Given the lack of spatial co-location of the burst source and this star, it is unlikely that these two objects are of a common origin.
We also queried NASA’s The High Energy Astrophysics Science Archive Research Center (HEASARC) within a radius of 2′′ to search for any high energy counterpart. The query returned an XMM Newton Source at J2000 RA and Dec 18:18:34.639, –15:31:35.1 with an error radius of 1.316′′. It has a combined EPIC band 8 flux of . For a distance of 10.9 kpc, this would become an X-ray luminosity of . The position of this X-ray source is not consistent with the realfast position within the formal error, but is separated at a distance within 2 of the positional error.
@@ 5.1.2 Supernova Remnant and Pulsar Wind Nebula
It does not appear that our burst source is a pulsar associated with this PWN or supernova event, based simply on its large offset; if it is associated with the PWN, the pulsar (or Fermi emission) should be directly coincident with that emission. Neither the Fermi or realfast error regions are coincident with the PWN emission. The realfast position sits at a distance of 6.1′ from the central position of the PWN. At the distance of 9.3 kpc indicated by Su et al. (2017), and their modelled supernova age of 11 kyr, the pulsar would have had to travel at a velocity of 1400 km s-1 to reach its current position (this minimum assumes the trajectory is perfectly orthogonal to our line-of-sight). This is much larger than typical pulsar birth velocities in the Milky Way (Lorimer et al., 1997). While there is some precedent for such a configuration (Schinzel et al., 2019), the lack of morphological evidence of such a high velocity (trail or bow shock) in the VLA and archival data, in addition to the central PWN, argue against that scenario.
@@ 5.1.3 Statistical association of the Fermi error regions
First, it is clear that the Fermi 3FGL position is slightly different from that of 4FGL. Both 95% Fermi error regions are consistent with the realfast bursts. The 4FGL position has a much smaller error region, and we characterize a conservative likelihood of chance coincidence for this occurring by answering the question, “given a random realfast position, what’s the chance it would have landed within a random 4FGL error region of this size?” We do this by estimating the density of Fermi sources (number of sources per square degree) near the Galactic center region, then multiplying this by the angular coverage of the 4FGL region in Figure 8. In the 4FGL catalog, we find 559 sources within the region covering in Galactic longitude and in Galactic latitude around the Galactic center. This tells us the source density per sky area is around 0.7 sources per square degree. Multiplying that by the area of the 4FGL ellipse (minor axis 144.36′′, major axis 194.76′′), we find a probability of 0.0048 that the realfast position would overlap the 4FGL error region by chance. This number would be less if we considered only unidentified Fermi sources, and took into account the additional confidence afforded by the 3FGL error region; thus, we consider the co-location, and likely association of the Fermi source with this burst emitter, to be relatively secure.
If the source were associated with a -ray pulsar then following equation 24 of the Smith et al. (2023), , using the DM distance of 10.9 kpc and a phase-averaged integral energy flux in the 0.1 to 300 GeV energy band = erg cm-2 s-1, which yields a -ray luminosity of erg s-1 with the beaming fraction, . If the nature of the detected pulses were from a pulsar-like object, this would put it into the category of young pulsars.
@@ 5.2.1 Magnetars
Magnetars are a class of rotating neutron stars with very high magnetic fields, and belong to the younger class of neutron stars. They can exhibit a rotation period of 1-10 s and very high spin down up to . They typically show time-variable flux densities, high linear polarization, profile changes, and flat spectral index (Levin et al., 2012). Some magnetars are known to show an inverted spectrum (Camilo et al., 2007a; Levin et al., 2012). Some magnetars also show circular polarization (Levin et al., 2012; Kramer et al., 2007). These are also known -ray and X-ray emitters in keV and MeV ranges. Their decaying magnetic field powers their emission, and they are capable of emitting radio bursts of energies similar to extragalactic FRBs, up to (Bochenek et al., 2020; CHIME/FRB Collaboration et al., 2020).
The high polarization fraction, variable fluxes, and energies of J1818–1531 are consistent with a magnetar origin. Magnetars are also known to have a large duty cycle up to 50% (Levin et al., 2012), which can lead to a non-detection of a period. The gradual change or flatness of PA is also commonly seen in magnetars (Camilo et al., 2007b).
Magnetars have not been detected in GeV energies, except for one extragalactic magnetar giant flare detected by Fermi-LAT (Fermi-LAT Collaboration et al., 2021). However, that flare was transient (fading on a timescale of less than 300 seconds). The high energy magnetar thermal emission typically arises in the X-ray band and is modelled by a blackbody in the soft X-ray band and a power law component in the hard X-rays. Takata et al. (2013) predicts non-thermal -ray luminosity from magnetars of the order in the GeV energies and suggests that the GeV photons can only escape the magnetar surface if the corresponding X-ray luminosity is . If the XMM Newton source is associated with J1818–1531, then the X-ray and -ray energies are consistent with this model and this will be the first detection of GeV energies from a Galactic magnetar. However, the offset between the X-ray and the radio sources currently disfavors this association.